OpenStudy (anonymous):

Please help... what is the equation of all the circles with fixed radius r and tangent to the x-axis...

OpenStudy (anonymous):

center of the circle will be (x,y)=(a,-r)

OpenStudy (anonymous):

can you give me an equation..:))

OpenStudy (anonymous):

or (x,y)=(a,r) then equation will be

OpenStudy (anonymous):

\[(x-a)^2+(y \pm r)^2=r^2\]

OpenStudy (anonymous):

why is it+-?

OpenStudy (anonymous):

that means + or - and u have 2 equation \[(x-a)^2+(y-r)^2=r^2 \\ \\ (x-a)^2+(y+r)^2=r^2\]

OpenStudy (anonymous):

ok.. thank you very much:))

OpenStudy (anonymous):

yw